# Bayesian characterization of the young open cluster NGC 6383 using HDBSCAN and Gaia DR3

### v1

@@ 2 Methodology
We acquired data from the Gaia third Data Release (DR3), executing a cone search of $25\leavevmode\nobreak\ \mathrm{arcmin}$ radius on the Gaia archive, yielding $71847$ sources. Initial selection filtered sources to distances between $500\leavevmode\nobreak\ \mathrm{pc}$ and $2500\leavevmode\nobreak\ \mathrm{pc}$—equivalent to parallax ranges between $0.4\leavevmode\nobreak\ \mathrm{mas}$ and $2\leavevmode\nobreak\ \mathrm{mas}$—based on limits established in prior research, resulting in $24476$ sources.
@@
To enhance data reliability, we applied astrometric fidelity parameters from Rybizki et al. (2022). This parameter, derived from a neural network analysis of $17$ Gaia catalog metrics, assesses the astrometric solution’s trustworthiness. Sources with an astrometric fidelity above $0.5$ were retained, narrowing the selection to $16857$ sources.
Subsequent refinement ensured inclusion of only sources with comprehensive parameters, finalizing the dataset at $16640$ sources. Systematic parallax offsets identified in Lindegren et al. (2021) were corrected using the Gaiadr3_zeropoint package. Moreover, addressing the bias in proper motion for bright sources ($\mathrm{G}<13\leavevmode\nobreak\ \mathrm{mag}$), as discussed by Cantat-Gaudin & Brandt (2021), we applied a magnitude-based correction for sources with $\mathrm{G}=11-13\leavevmode\nobreak\ \mathrm{mag}$, compensating for up to $80\leavevmode\nobreak\ \mathrm{\mu as\,yr^{-1}}$ discrepancy between the frames of reference for bright and faint sources.
@@ 2.1 COSMIC
COSMIC or Characterization Of Star clusters using Machine learning Inference and Clustering, developed by Lucas Pulgar-Escobar et al. (in prep), is a suite of functions designed for analyzing open clusters. Utilizing unsupervised machine learning algorithms, COSMIC processes extensive datasets, such as those from Gaia, to identify fundamental parameters of open clusters through clustering techniques and Bayesian estimation. As an open-source program, it is developed in Python 3.11 and integrates PyMC 5.10222www.pymc.io, a Python library specialized in Bayesian analysis.
@@ 2.1.1 Membership determination
Figure 3: Left: Gaia G-band magnitude ($G_{\mathrm{mag}}$) vs. parallax ($\varpi$), its noticeable that sources with higher magnitudes, hence fainter, exhibit larger parallax errors, rendering them less reliable for certain analyses. Blue dots are sources with fractional parallax less than $0.1$, and red ones are sources with values over $0.1$, being excluded for the distance estimation. Mean parallax is indicated by the green dashed line. Right: Sample of the inverse values of the parallaxes as distances.
After narrowing the data to a final subset of $16640$ sources, HDBSCAN was applied [^3] The algorithm’s hyperparameters included an Euclidean distance metric and a minimum cluster size of $43$., focusing on proper motions as the primary clustering parameters.
Finally, we used the AstroPy Sigma Clipping (Astropy Collaboration et al., 2022) utility to perform a $2\sigma$ clipping method around the median in order to obtain the final likely members of the cluster.
@@ 2.1.2 Parallax and Distance Estimation
Figure 4: Radial density profile of the cluster. Obtained King profiles are indicated with shaded blue zone, and the magenta line indicated the median of the profiles, while the blue dots correspond to the calculated density, structural parameters are indicated by dashed and dotted lines.
@@ 2.1.3 Proper motions
A normal prior distribution was assigned to the mean proper motions $\left[\overline{\mu_{\alpha}},\overline{\mu_{\delta}}\right]$, based on frequentist means and standard deviations. The standard deviations are modeled using a HalfNormal distribution, while the correlation coefficient between the proper motions is uniformly distributed over the interval $[-1,1]$.
@@ 2.1.5 Radial density profile
To determine structural parameters of the cluster, we utilized the King (1962) density profile.
The numerical density $\rho$, was computed by dividing the cluster area into concentric annuli, each containing an equal number of stars. The number of annuli, $K$, adheres to the equiprobable bin rule ($K=2n^{2/5}$), where $n$ is the star count within the cluster. Fitting King’s model to the density data allows for estimating the structural parameters’ posterior distributions.
Priors for the model parameters were set as follows: $b\sim\mathcal{U}(0,2\rho_{\mathrm{min}})$, $k\sim\mathcal{U}(0,2\rho_{\mathrm{max}})$, $R_{c}\sim\mathcal{U}(0,0.8R_{t})$, and $R_{t}\sim\mathcal{U}(R_{c},1.5T_{\mathrm{max}})$, where $T_{\mathrm{max}}$ represents the highest value among tidal potential radius, Hill radius, gravitational bound radius, and the maximum observed cluster member distance. This ensures the chosen parameter space for $R_{c}$ and $R_{t}$ is within astrophysical valid limits.
@@ 2.2 AsTeCA
AsTeCA, or Automated Stellar Cluster Analysis (Perren et al., 2015), is a suite of tools designed to automate the standard tests applied to stellar clusters to determine their basic parameters. In order to obtain accurate estimates for a cluster’s metallicity, age and extinction values, we use AsTeCA’s isochrone fitting process. We used the PARSEC v1.2S isochrones (Bressan et al., 2012; Tang et al., 2014), with the Gaia EDR3 photometric system and Kroupa (2001, 2002) canonical two-part-power law IMF corrected for unresolved binaries, with a logarithmic age range between $6.0$ and $7.9$. We select use the median solution, and all the parameters prior are set as uniform distributions.
@@
Distance $(\varpi)$ $1.090\pm 0.050$ $\mathrm{kpc}$
@@
Parallax $(\varpi)$ $0.919\pm 0.042$ $\mathrm{mas}$
Number of members 304 $\mathrm{stars}$
@@
Core Radius $(R_{c})$ $1.402\pm 0.1593$ $\mathrm{arcmin}$
Background $(b)$ $0.0390\pm 0.020$ $\mathrm{stars\,arcmin^{-2}}$
@@
Tidal Radius $(R_{t})$ $29.331\pm 8.507$ $\mathrm{arcmin}$
Center Density $(k)$ $10.476\pm 0.965$ $\mathrm{stars\,arcmin^{-2}}$
Cluster Center R.A. $263.672\pm 0.004$ $\mathrm{deg}$
Cluster Center DEC. $-32.577\pm 0.004$ $\mathrm{deg}$
Proper Motion R.A. $2.534\pm 0.009$ $\mathrm{mas\,yr^{-1}}$
Proper Motion DEC. $-1.711\pm 0.007$ $\mathrm{mas\,yr^{-1}}$
@@ 3 Preliminary Results
Table 1 presents the results of our study, focusing on parameters derived from sources with a membership probability exceeding 60 percent. Fig.1 illustrates the cluster’s center, while proper motions is depicted in Fig.2. Additionally, Fig.3 provides insights into the distance and parallax estimation, and Fig.4 showcases the cluster’s structural parameters alongside the fitted King Profile. Furthermore, Fig.5 displays the best-fit isochrone, and the CMD of NGC 6383.
@@ 4 Conclusion and future work
Our research applied Bayesian inference and HDBSCAN clustering algorithm to Gaia DR3 data to advance the characterization of NGC 6383. We have refined key parameters such as distance, proper motion, and age.
Future work should aim at expanding membership analysis to achieve a more comprehensive cluster census, and integrating radial velocity data to explore cluster dynamics further. Also, the use of a larger search cone to find possible stars in the tidal region of the cluster. Additionally, incorporating multi-wavelength data could enhance the accuracy of determining stellar properties. Finally, a focused investigation into the binary and multiple systems within NGC 6383 could offer valuable perspectives on its star formation history and initial mass function, for example more in detail membership of the central binary HD 159176. {acknowledgement} We gratefully acknowledge support by the ANID BASAL project FB210003.

### v2

@@ 2 Methodology
We acquired data from the Gaia third Data Release (DR3), executing a cone search of $25\leavevmode\nobreak\ \mathrm{arcmin}$ radius on the Gaia archive, yielding $71847$ sources. Initial selection filtered sources to distances between $500\leavevmode\nobreak\ \mathrm{pc}$ and $2500\leavevmode\nobreak\ \mathrm{pc}$—equivalent to parallax ranges between $0.4\leavevmode\nobreak\ \mathrm{mas}$ and $2\leavevmode\nobreak\ \mathrm{mas}$—based on limits established in prior research, resulting in $29092$ sources.
@@
To enhance data reliability, we applied astrometric fidelity parameters from Rybizki et al. (2022). This parameter, derived from a neural network analysis of $17$ Gaia catalog metrics, assesses the astrometric solution’s trustworthiness. Sources with an astrometric fidelity above $0.5$ were retained, narrowing the selection to $20215$ sources.
Subsequent refinement ensured inclusion of only sources with comprehensive parameters, finalizing the dataset at $19964$ sources. Systematic parallax offsets identified in Lindegren et al. (2021) were corrected using the Gaiadr3_zeropoint package. Moreover, addressing the bias in proper motion for bright sources ($\mathrm{G}<13\leavevmode\nobreak\ \mathrm{mag}$), as discussed by Cantat-Gaudin & Brandt (2021), we applied a magnitude-based correction for sources with $\mathrm{G}=11-13\leavevmode\nobreak\ \mathrm{mag}$, compensating for up to $80\leavevmode\nobreak\ \mathrm{\mu as\,yr^{-1}}$ discrepancy between the frames of reference for bright and faint sources.
@@ 2.1 COSMIC
Characterization Of Star clusters using Machine learning Inference and Clustering (COSMIC) developed by Lucas Pulgar-Escobar et al. (in prep), is a suite of functions designed for analyzing open clusters. Utilizing unsupervised machine learning algorithms, COSMIC processes extensive datasets, such as those from Gaia, to identify fundamental parameters of open clusters through clustering techniques and Bayesian estimation. As an open-source program, it is developed in Python 3.11 and integrates PyMC 5.10222www.pymc.io, a Python library specialized in Bayesian analysis.
@@ 2.1.1 Membership determination
Figure 3: Left panel: Gaia G-band magnitude ($G_{\mathrm{mag}}$) vs. parallax ($\varpi$), it’s noticeable that fainter sources exhibit larger parallax errors, rendering them less reliable. Blue dots are sources with fractional parallax less than $0.1$, and red ones are sources with values over $0.1$, being excluded for the distance estimation. The green dashed line indicates mean parallax. Right panel: Histogram of the inverse values of the parallaxes as distances, with the same criteria as the left panel.
HDBSCAN was applied [^3] The algorithm’s hyperparameters included an Euclidean distance metric and a minimum cluster size of $43$. to the $19964$ sources using the proper motions as the clustering parameters, resulting in 544 sources with a probability over $0.5$.
We applied the AstroPy Sigma Clipping utility (Astropy Collaboration et al., 2022) for a $2\sigma$ clipping around the median, yielding 399 probable members for parameters inference.
@@ 2.1.3 Proper motions
A normal prior distribution was assigned to the mean proper motions, based on frequentist means and standard deviations. The standard deviations are modeled using a HalfNormal distribution, while the correlation coefficient between the proper motions follows a $\mathcal{U}(-1,1)$.
Figure 4: Radial density profile of NGC 6383, with the blue dots representing the stellar density. The shaded blue zone outlines the deviation of the King’s profiles, and the magenta line represents the median profile. The core radius $R_{c}$ is shown as a red dashed line, the tidal radius $R_{t}$ as a red dotted line, and the background density $b$ as a green dashed line. The logarithmic axes denote the radius $r$ in arcminutes and the stellar density $\rho$ per arcminute squared.
@@ 2.1.5 Radial density profile
To determine structural parameters of the cluster, we utilized the King (1962) density profile. The numerical density $\rho$, was computed by dividing the cluster area into concentric annuli, each containing an equal number of stars. The number of annuli, $K$, adheres to the equiprobable bin rule ($K=2n^{2/5}$), where $n$ is the star count within the cluster.
Model priors were set as follows: background density $b\sim\mathcal{U}(0,2\rho_{\mathrm{min}})$, scale factor $k\sim\mathcal{U}(0,2\rho_{\mathrm{max}})$, core radius $R_{c}\sim\mathcal{U}(0,0.8R_{t})$, and tidal radius $R_{t}\sim\mathcal{U}(R_{c},1.5T_{\mathrm{max}})$. $T_{\mathrm{max}}$ is the maximum value among tidal potential radius, Hill radius, gravitational bound radius, and the maximum observed cluster member distance. This ensures the chosen parameter space for $R_{c}$ and $R_{t}$ is within astrophysical valid limits.
@@ 2.2 AsTeCA
AsTeCA, or Automated Stellar Cluster Analysis (Perren et al., 2015), is a suite of tools designed to automate the standard tests applied to stellar clusters to determine their basic parameters. In order to obtain accurate estimates for a cluster’s metallicity, age and extinction values, we use AsTeCA’s isochrone fitting process. We used the PARSEC v1.2S isochrones (Bressan et al., 2012; Tang et al., 2014), with the Gaia EDR3 photometric system and Kroupa (2001, 2002) canonical two-part-power law IMF corrected for unresolved binaries, with a logarithmic age range between $6.0$ and $7.9$. All parameter priors are set as uniform distributions, and the median values of the resulting posterior distributions are informed.
@@
Distance $(\varpi)$ $1.099\pm 0.037$ $\mathrm{kpc}$
@@
Parallax $(\varpi)$ $0.917\pm 0.060$ $\mathrm{mas}$
Number of members $266$ $\mathrm{stars}$
@@
Core Radius $(R_{c})$ $1.205\pm 0.126$ $\mathrm{arcmin}$
Background $(b)$ $0.035\pm 0.017$ $\mathrm{stars\,arcmin^{-2}}$
@@
Tidal Radius $(R_{t})$ $29.732\pm 7.694$ $\mathrm{arcmin}$
Center Density $(k)$ $10.508\pm 0.795$ $\mathrm{stars\,arcmin^{-2}}$
Cluster Center R.A. $263.673\pm 0.004$ $\mathrm{deg}$
Cluster Center DEC. $-32.580\pm 0.004$ $\mathrm{deg}$
Proper Motion R.A. $2.541\pm 0.007$ $\mathrm{mas\,yr^{-1}}$
Proper Motion DEC. $-1.714\pm 0.006$ $\mathrm{mas\,yr^{-1}}$
@@ 3 Preliminary Results
The informed results and figures are obtained using sources with a membership probability exceeding 60 percent, ensuring a high level of confidence in the analysis. Table 1 presents a summary of the results, highlighting key parameters of NGC 6383. The cluster’s center is accurately defined in Fig.1, while Fig.2 illustrates the proper motions. Insights into distance and parallax estimation are provided in Fig.3, and Fig.4 showcases the cluster’s structural parameters alongside the fitted King’s Profile. Additionally, Fig.5 displays the CMD of NGC 6383 alongside the best-fit isochrone.
@@ 4 Conclusion and future work
Our study employed Bayesian methods and HDBSCAN to refine parameters like distance, proper motion, and age for NGC 6383 using Gaia DR3.
Future efforts will broaden member analysis for a complete census and incorporate radial velocity for dynamic insights. Also, the use of a larger search cone to find possible stars in the tidal region of the cluster. Additionally, incorporating multi-wavelength data could enhance the accuracy of determining stellar properties. Finally, a focused investigation into the binary and multiple systems within NGC 6383 like the central binary HD 159176 could offer valuable information about its star formation history. {acknowledgement} We gratefully acknowledge support by the ANID BASAL project FB210003.